OpenStudy (anonymous):

Is this right? I had to find the interior and arc length of a pizza I created with a diameter of 14 inches. Solved for the Pizza with 8 slices C= pi x d or C= pi x 14= 44 Circumference is 44 inches 44/8= 5.5 inches divided into 8 pieces 360/8= 45 Each angle is 45 degrees 45/360= 1/8 45/360(pd) 45/360 (px14) 0.125(14)(p) 1.75(pi)= 5.49 which rounds to 5.5 The interior angle is 45 degrees and the arc length is 5.5inches. Area for 8 sliced pizza Area of whole circle divided by 8 A= pi x r2/8 A= pi x 142/8 or = pi x 196/8 A= pi x 24.5= 77.07 in2

OpenStudy (anonymous):

I meatnt to say "interior angle"

OpenStudy (anonymous):

yes that looks correct :)

OpenStudy (anonymous):

Thanks! anyone else care to comment?

OpenStudy (callisto):

C= pi x d or C= pi x 14= 44 Circumference is 44 inches 44/8= 5.5 inches divided into 8 pieces 360/8= 45 Each angle is 45 degrees ^ correct

OpenStudy (anonymous):

gotcha. Thanks!

OpenStudy (callisto):

45/360= 1/8 45/360(pd) 45/360 (px14) 0.125(14)(p) 1.75(pi)= 5.49 which rounds to 5.5 ^ no need you found that before

OpenStudy (callisto):

A= pi x r2/8 A= pi x 142/8 or = pi x 196/8 A= pi x 24.5= 77.07 in2 ^ incorrect

OpenStudy (callisto):

14 is the diameter radius = diameter/2 = 14/2 =7

OpenStudy (anonymous):

Ohhh crud I see what you mean now

OpenStudy (callisto):

area of each slice\[\pi r ^{2} /8\] = π(7)^2 /8 = 19.25

OpenStudy (anonymous):

wouldn't it be 19.24? there is only a 2 after the 4

OpenStudy (callisto):

22/7 x 7x7 /8 = 22x7/8 = 19.25

OpenStudy (anonymous):

ohhh. I/m using he number pi not the fraction